OpenStudy (anonymous):

A certain tennis player makes a successful serve 64% of the time. Assume that each serve is independant of the others. If she serves 9 times, what is the probability she a) serves all 9 in? b)serves exactly 8 in?c) serves at least 7 in? d) serves no more than 7 in?

OpenStudy (campbell_st):

well its binomial probability p = 0.64 serve success q = 0.36 serve failure so the probability of k successes from n trials is \[P(k ..successes.. from ..9 ..trials) = ^9C_{k} (0.64)^{k}(0.36)^{9 - k}\] so for the 1st question k = 9 2nd k = 8 3rd, at least 7 is k = 7 + k = 8 + k = 9 4th 1 - (k = 8 + k = 9) hope this helps